# Unified Source-Free Domain Adaptation

### v1

@@ 3.5 Model Training
We train LCFD in an alternating manner. In each iteration, we first optimize prompt context $\boldsymbol{v}_{e}$ with ${\cal L}_{{\rm{EC}}}$ whilst the target model is frozen (i.e., phase-1), followed by training the target model by ${\cal L}_{{\rm{IC}}}$ (i.e., phase-2) while the prompt context is frozen. The ViL model is always frozen throughout. More details for training is presented in Alg. 1.
TABLE II: Closed-set SFDA results (%) on Office-Home and VisDA. SF and M means source-free and multimodal, respectively.
Method Venue SF M Office-Home VisDA
@@
CLIP-RN ICML21 51.9 81.5 82.5 72.5 81.5 82.5 72.5 51.9 82.5 72.5 51.9 81.5 72.1 83.7
CLIP-B32 ICML21 59.8 84.3 85.5 74.6 84.3 85.5 74.6 59.8 85.5 74.6 59.8 84.3 76.1 82.9
DAPL-RN TNNLS23 54.1 84.3 84.8 74.4 83.7 85.0 74.5 54.6 84.8 75.2 54.7 83.8 74.5 86.9
PADCLIP-RN ICCV23 57.5 84.0 83.8 77.8 85.5 84.7 76.3 59.2 85.4 78.1 60.2 86.7 76.6 88.5
ADCLIP-RN ICCV23W 55.4 85.2 85.6 76.1 85.8 86.2 76.7 56.1 85.4 76.8 56.1 85.5 75.9 87.7
SHOT ICML20 55.0 78.7 81.3 69.1 78.9 79.1 68.2 53.6 81.6 73.5 59.4 83.5 71.8 82.9
GKD IROS21 56.5 78.3 82.2 69.2 80.4 78.7 67.4 55.4 82.6 74.3 60.3 84.2 72.5 83.0
NRC NIPS21 57.2 79.3 81.3 68.9 80.6 80.2 66.6 57.3 82.0 71.0 57.9 84.9 72.3 85.9
AaD NIPS22 59.3 79.3 82.1 68.9 79.8 79.5 67.2 57.4 83.1 72.1 58.5 85.4 72.7 88.0
AdaCon CVPR22 47.2 75.1 75.5 60.7 73.3 73.2 60.2 45.2 76.6 65.6 48.3 79.1 65.0 86.8
CoWA ICML22 57.3 79.3 81.0 69.3 77.9 79.6 68.1 56.4 82.6 72.9 61.3 83.7 72.4 86.9
PLUE CVPR23 49.1 73.5 78.2 62.9 73.5 74.5 62.2 48.3 78.6 68.6 51.8 81.5 66.9 88.3
TPDS IJCV23 59.3 80.3 82.1 70.6 79.4 80.9 69.8 56.8 82.1 74.5 61.2 85.3 73.5 87.6
LCFD-C-RN 60.1 85.6 86.2 77.2 86.0 86.3 76.6 61.0 86.5 77.5 61.4 86.2 77.6 89.3
LCFD-C-B32 72.3 89.8 89.9 81.1 90.3 89.5 80.1 71.5 89.8 81.8 72.7 90.4 83.3 89.3
TABLE III: Closed-set SFDA results (%) on DomainNet-126. SF and M means source-free and multimodal, respectively.
@@
CLIP-RN ICML21 70.2 87.1 65.4 67.9 87.1 65.4 67.9 70.2 65.4 67.9 70.2 87.1 72.7
CLIP-B32 ICML21 73.5 85.7 71.2 74.7 85.7 71.2 74.7 73.5 71.2 74.7 73.5 85.7 76.3
DAPL-RN TNNLS23 72.4 87.6 65.9 72.7 87.6 65.6 73.2 72.4 66.2 73.8 72.9 87.8 74.8
ADCLIP-RN ICCV23W 71.7 88.1 66.0 73.2 86.9 65.2 73.6 73.0 68.4 72.3 74.2 89.3 75.2
SHOT ICML20 63.5 78.2 59.5 67.9 81.3 61.7 67.7 67.6 57.8 70.2 64.0 78.0 68.1
GKD IROS21 61.4 77.4 60.3 69.6 81.4 63.2 68.3 68.4 59.5 71.5 65.2 77.6 68.7
NRC NIPS21 62.6 77.1 58.3 62.9 81.3 60.7 64.7 69.4 58.7 69.4 65.8 78.7 67.5
AdaCon CVPR22 60.8 74.8 55.9 62.2 78.3 58.2 63.1 68.1 55.6 67.1 66.0 75.4 65.4
CoWA ICML22 64.6 80.6 60.6 66.2 79.8 60.8 69.0 67.2 60.0 69.0 65.8 79.9 68.6
PLUE CVPR23 59.8 74.0 56.0 61.6 78.5 57.9 61.6 65.9 53.8 67.5 64.3 76.0 64.7
TPDS IJCV23 62.9 77.1 59.8 65.6 79.0 61.5 66.4 67.0 58.2 68.6 64.3 75.3 67.1
LCFD-C-RN 75.4 88.2 72.0 75.8 88.3 72.1 76.1 75.6 71.2 77.6 75.9 88.2 78.0
LCFD-C-B32 77.2 88.0 75.2 78.8 88.2 75.8 79.1 77.8 74.9 79.9 77.4 88.0 80.0
4 Experiments
4.1 Implementation Details
Source models. Following [44], for source-free out-of-distribution generalization (SF-OODG) we adopt the Pytorch built-in ResNet50 model (pre-trained on ImageNet) as the source model, taking ImageNet variations as evaluation datasets. For all the other settings, the source model is trained on the source domain in a supervised manner, same as [15, 18, 20].
Networks. Our LCFD involves two networks: the ViL model and the target model. In practice, we choose CLIP [11] as the ViL model. The target model’s structure is the same as the source model. Following [15, 19], a target model consists of a feature extractor and a classifier (a FC layer). We adapt the feature extractor, which pre-trained on ImageNet. For fair comparison, we use ResNet101 as the backbone for VisDA and ResNet50 for the others.
Training. There are three parameters with LCFD: $\alpha$ in $L_{\rm{EC}}$ (Eq. (13)), $\sigma$ and $\tau$ in $L_{\rm{IC}}$ (Eq. (14)). For all settings, we adopt the same configuration $(\alpha,\sigma,\tau)=(0.003,0.4,1.0)$. Here, since the value of $\mathcal{{L}}_{\rm{VMI}}$ is large, $\alpha$ is set to the small one of 0.003. For model training, we use the batch size of 64, SGD optimizer with momentum 0.9 and 15 epochs on all datasets. The prompt template for initialization is ’a photo of a [CLASS].’ [11].
@@ 4.2 Closed-Set SFDA
Data sets. In this part, we evaluate LCFD on three challenging benchmarks for domain adaptation. Office-Home [45] is a middle-scale dataset. It contains 15k images belonging to 65 categories from working or family environments, being divided into four domains, i.e., Artistic images, Clip Art, Product images, and Real-word images. VisDA [46] is a large-scale dataset. It includes 12 types of Synthetic to Real transfer recognition tasks. The source domain contains 152k synthetic images, whilst the target domain has 55k real object images from Microsoft COCO. DomainNet-126 [47] is another large-scale dataset. As a subset of DomainNet containing 600k images of 345 classes from 6 domains of different image styles, this dataset has 145k images from 126 classes, sampled from 4 domains, Clipart, Painting, Real, Sketch, as [47] identifies severe noisy labels in the dataset.
TABLE IV: One step adapting results (%) on Office-Home in Generalized SFDA setting. S, T, H are the results on the source and target domains, and harmonic mean accuracy, respectively; the bracket values in red are the gap between S and T; WAD means With Anti-forgetting Design.
@@
S T (S-T gap) H
Source 98.1 59.2 (38.9) 73.1
SHOT ICML20 84.2 71.8 (12.4) 77.1
GKD IROS21 86.8 72.5 (14.3) 78.7
NRC NIPS21 91.3 72.3 (19.0) 80.4
AdaCon CVPR22 88.2 65.0 (23.2) 74.4
CoWA ICML22 91.8 72.4 (19.4) 80.6
PLUE CVPR23 96.3 66.9 (29.4) 78.4
TPDS IJCV23 83.8 73.5 (10.3) 78.0
GDA ICCV21 80.0 70.2 (9.8) 74.4
PSAT-ViT TMM23 86.4 83.6 (2.8) 85.0
LCFD-C-RN 85.0 77.6 (7.4) 80.7
LCFD-C-B32 86.3 83.3 (3.0) 84.5
TABLE V: Continual adaptation results (%) on Office-Home under Generalized SFDA setting, the first column of each sub-table indicates the adaptation sequence. $\downarrow$ means the average accuracy drop of a test domain on the adaptation path compared with the performance when the domain is first seen.
ResNet50 (RN)
@@
Competitors. To evaluate LCFD, we select 14 state-of-the-art comparisons divided into three groups. (i) The first is the source model and CLIP that lay the basis for LCFD. (ii) The second is the UDA method DAPL [48], PADCLIP [49], ADCLIP [50] that introduces prompt learning to boost the cross-domain transfer. (iii) The third includes 9 SFDA methods: SHOT [15], GKD [25], NRC [18], SFDE [6], AdaCon [51], CoWA [23], PLUE [17], TPDS [52]. According to idea, SHOT, GKD, CoWA are pseudo-labels-based methods; NRC, PLUE, AdaCon, SFDE exploit the data geometry, e.g., the nearest neighbor, source prototype; TPDS predicts the target probability distribution using error control.
For comprehensive comparisons, we implement LCFD in two variants: (i) LCFD-C-RN (base) and (ii) LCFD-C-B32 (premium). The key distinction lies in the backbone of the CLIP image-encoder. Specifically, for LCFD-C-RN, ResNet101 is employed on the VisDA dataset, while ResNet50 is used on the Office-Home and DomainNet-126 datasets. On the other hand, LCFD-C-B32 adopts ViT-B/32 [53] as the backbone across all datasets. Similarly, all comparison methods using CLIP are marked with the same meaning for a convenience observation.
@@
On the other hand, LCFD-C-RN defeats CLIP-RN on all tasks, achieving the improvement of 4.5%, 5.6% and 5.3% on Office-Home, VisDA and DomainNet-126 in average, respectively. Similarly, compared with CLIP-B32, the corresponding improvement of LCFD-C-B32 are changed to 7.2%, 6.4% and 3.7%. This indicates the effect of the prompt learning to discover the external causal elements.
4.3 Generalized SFDA
Data sets. This setting evaluates LCFD on the Office-Home, following previous generalized SFDA methods [3, 8].
Evaluation protocol. Unlike closed-set SFDA, the generalized SFDA problem highlights the anti-forgetting ability on the seen source domain. In terms of evaluation rule, the same as [3], we adopt the harmonic mean accuracy that is computed by ${H}=({2*A_{s}*A_{t}})/({A_{s}+A_{t}})$ where $A_{s}$ and $A_{t}$ are the accuracy on the source domain and the target domain, respectively. Note that the $A_{s}$ is computed based on the source-testing set. The same to [3, 8], on the source domain, the ratio of training and testing sets is 9:1.
Competitors. As listed in Tab. IV, two generalized SFDA methods GDA [3], PSAT [8] are additionally selected as the comparisons. In GDA, domain attention was integrated with a local clustering-based self-supervised learning method. Essentially, the capacity for the old tasks is kept by encouraging the model parameters closed to the original values. Unlike GDA, PSAT enforces the adaptation process to remember the source domain by imposing source guidance, building a target domain-centric anti-forgetting mechanism.
One step adapting comparison. Tab. IV reports the results on Office-Home, including H-accuracy, S-accuracy and T-accuracy. In the average H-accuracy aspect, LCFD-C-B32 is only defeated by generalized SFDA method PSAT (having particular design of anti-forgetting mechanism and based on ViT backbone) with a gap of 0.5%. Adopting the weak version LCFD-C-RN still outperforms other methods except for the best PSAT. Although T-accuracy of NRC, CoWA closes LCFD-C-RN due to the high S-accuracy (91%), the S-T gap of them (19%) is much larger than LCFD-C-RN (7.4%). These results imply a better balance of LCFD as one-step adapting.
[…]
4.4 Open-Set & Patial-Set SFDA
Data sets. In this setting, we adopt Office-Home as the evaluation benchmark, as done in [15, 9].
TABLE VI: Partial-set SFDA and open-set SFDA results on the Office-Home dataset (%).
@@ 4.4 Open-Set & Patial-Set SFDA
SHOT ICML20 79.3 SHOT ICML20 72.8
HCL NIPS21 79.6 HCL NIPS21 72.6
CoWA ICML22 83.2 CoWA ICML22 73.2
AaD NIPS22 79.7 AaD NIPS22 71.8
CRS CVPR23 80.6 CRS CVPR23 73.2
LCFD-C-RN 82.9 LCFD-C-RN 79.6
LCFD-C-B32 85.8 LCFD-C-B32 83.4
Competitors. Besides methods Source, CLIP, SHOT, CoWA from the closed-set setting, this part adds three new comparisons HCL [54], AaD [55] and CRS [9] that all follow the contrastive learning framework.
Comparison results. As the comparison results on Office-Home presented in Tab. VI, in average accuracy, LCFD-C-B32 surpasses not only all comparisons with an improvement of 2.6% and 10.2% compared with previous work CoWA (partial-set) and CRS (open-set). As for LCFD-C-RN, only CoWA beat it in the partial-set setting with a tiny leading of 0.3%. The quantitative results indicate that LCFD is still competitive on the partial-set and open-set settings.

### v4

@@ 3.5 Model training
We train CausalDA in an alternating manner. In each iteration, we first optimize prompt context $\boldsymbol{v}_{e}$ with ${\cal L}_{{\rm{EC}}}$ whilst the target model is frozen (i.e., phase-1), followed by training the target model by ${\cal L}_{{\rm{IC}}}$ (i.e., phase-2) while the prompt context is frozen. The ViL model is always frozen throughout. More details for training are presented in Alg. 1.
4 Experiments
4.1 Implementation details
Source models. Following [70], for source-free out-of-distribution generalization (SF-OODG) we adopt the Pytorch built-in ResNet50 model (pre-trained on ImageNet) as the source model, taking ImageNet variations as evaluation datasets. For other settings, the source model is trained on the source domain in a supervised manner, same as [31, 54].
Networks. Our CausalDA involves two parts: the ViL model and the target model. We choose CLIP [42] as the ViL model. The target model’s structure is the same as the source model. Following [31, 65], a target model consists of a feature extractor and a classifier (an FC layer). We adapt the feature extractor, which is pre-trained on ImageNet. For fair comparison, we use ResNet101 as the backbone for VisDA and ResNet50 for the others.
Training. There are three parameters with CausalDA: $\alpha$ in $L_{\rm{EC}}$ (Eq. (14)), $\tau$ in $L_{\rm{UN}}$ (Eq. (16)), and $\sigma$ in $L_{\rm{IC}}$ (Eq. (17)). For all settings, we adopt the same configuration $(\alpha,\sigma,\tau)=(0.003,0.4,1.0)$. Here, since the value of $\mathcal{{L}}_{\rm{VMI}}$ is large, $\alpha$ is set as small as 0.003. For model training, we use a batch size of 64, an SGD optimizer with momentum 0.9, and 15 epochs on all datasets. The learnable prompt template is initialized with ’a photo of a [CLS].’ [42], where the [CLS] means a specific target class name.
Unification evaluation protocol. In the SFDA context, unification does not necessarily imply that the proposed method must achieve the absolute best performance in every individual adaptation scenario. Instead, the essence of contributing a unified solution lies in its ability to provide stable and competitive performance across a wide range of settings in a unified manner, without the need of resorting to a diversity of scenario specific designs.
To evaluate the unification of methods, we adopt three harmonic metrics: (1) Overall mean $H_{all}$ that averages over all settings, (2) worst-case relative gap $H_{wrg}$ that is largest percentage shortfall of a method from the per-setting best across all settings, and (3) leave-one-setting-out averages $H_{loso}$ omits one setting at a time and averages the scores over the rest. For a specific method $m$ evaluated over a number of settings $\mathcal{S}$, the metrics can be computed by:
$\small\begin{split}&\mathrm{H}_{all}(m)=\frac{1}{|\mathcal{S}|}\sum_{s\in\mathcal{S}}x_{m,s};~\mathrm{H}_{wrg}(m)=\max_{s\in\mathcal{S}}\frac{b_{s}-x_{m,s}}{b_{s}};\\ &\mathrm{H}_{loso}(m,s^{\prime})=\frac{1}{|\mathcal{S}-1|}\sum_{s\in\mathcal{S}\setminus s^{\prime}}x_{m,s},\end{split}$ (18)
where $x_{m,s}$ specifies the score of method $m$ under a setting $s$; $b_{s}\!\!=\!\!\max_{m^{\prime}}x_{m^{\prime},s}$ is the per-setting best; $s^{\prime}$ is a omitted setting.
4.2 Indirect validation of causal factors
Validating the discovery of causal factors is significant yet non-trivial in causality learning, especially with latent causal factors as in this study. To that end, we adopt a general principle that there exists an intimate connection between invariance and causality useful for generalization [45, 1]. That being said, if a learned model enables varying environments generalization, it constitutes indirect evidence that the model has captured causal predictors (e.g., a causal explanation of an object, such as why it is a computer).
Concretely, we simulate a spectrum of viewing environments by imposing varying levels of Gaussian noise with gradually increased kernel sizes from 8 to 20. We compare the competitors without (SHOT [31]) and with (DIFO [53] and ProDe [52]) ViL knowledge (the image-encoder in CLIP adopts backbone ViT-B/32 [15]). As shown in Fig. 5, all competitors degrade clearly as the noise increases, whilst CausalDA levels off. These findings suggest the method captures invariant predictors that are compatible with causal explanations, though they do not prove causal recovery rigorously. To further examine if external and internal factors both are effective, we further derive a variant of CausalDA, named as CausalDAin, without the external causality discovery. Specifically, in CausalDAin, the pseudo-labels are directly derived from CLIP’s zero-shot predictions. This is consistent with external factors contributing to improved invariance, suggesting a possible complementarity with internal counterparts.
The validation above provides indirect yet preliminary evidence for causality capture. In the following, we present further supporting evidence through adaptation results across diverse SFDA settings.
[…]
TABLE III: Closed-set results (%) on Office-Home, VisDA. SF and ViL means source-free and ViL model-required, respectively.
Method Venue SF ViL Office-Home VisDA
@@ 4.2 Indirect validation of causal factors
CLIP-RN [42] ICML21 51.9 81.5 82.5 72.5 81.5 82.5 72.5 51.9 82.5 72.5 51.9 81.5 72.1 83.7
CLIP-B32 [42] ICML21 64.9 86.6 87.2 78.0 86.6 87.2 78.0 64.9 87.2 78.0 64.9 86.6 79.2 85.2
DAPL-RN [14] TNNLS23 54.1 84.3 84.8 74.4 83.7 85.0 74.5 54.6 84.8 75.2 54.7 83.8 74.5 86.9
PADCLIP-RN [25] ICCV23 57.5 84.0 83.8 77.8 85.5 84.7 76.3 59.2 85.4 78.1 60.2 86.7 76.6 88.5
ADCLIP-RN [47] ICCV23W 55.4 85.2 85.6 76.1 85.8 86.2 76.7 56.1 85.4 76.8 56.1 85.5 75.9 87.7
SHOT [31] ICML20 55.0 78.7 81.3 69.1 78.9 79.1 68.2 53.6 81.6 73.5 59.4 83.5 71.8 82.9
GKD [48] IROS21 56.5 78.3 82.2 69.2 80.4 78.7 67.4 55.4 82.6 74.3 60.3 84.2 72.5 83.0
NRC [64] NeurIPS21 57.2 79.3 81.3 68.9 80.6 80.2 66.6 57.3 82.0 71.0 57.9 84.9 72.3 85.9
AaD [63] NeurIPS22 59.3 79.3 82.1 68.9 79.8 79.5 67.2 57.4 83.1 72.1 58.5 85.4 72.7 88.0
AdaCon [5] CVPR22 47.2 75.1 75.5 60.7 73.3 73.2 60.2 45.2 76.6 65.6 48.3 79.1 65.0 86.8
CoWA [26] ICML22 57.3 79.3 81.0 69.3 77.9 79.6 68.1 56.4 82.6 72.9 61.3 83.7 72.4 86.9
PLUE [32] CVPR23 49.1 73.5 78.2 62.9 73.5 74.5 62.2 48.3 78.6 68.6 51.8 81.5 66.9 88.3
TPDS [49] IJCV24 59.3 80.3 82.1 70.6 79.4 80.9 69.8 56.8 82.1 74.5 61.2 85.3 73.5 87.6
SHOT-B32 [31] ICML20 64.7 84.0 86.4 79.4 87.3 85.4 78.1 65.8 87.6 81.6 65.0 88.5 79.5 85.2
GKD-B32 [48] IROS21 67.2 83.2 84.3 77.0 84.8 83.6 75.5 71.1 86.0 81.1 72.4 88.5 79.6 85.3
NRC-B32 [64] NeurIPS21 55.0 83.4 86.7 78.3 86.5 85.1 77.1 69.9 87.2 90.9 65.5 88.4 79.5 88.2
AaD-B32 [63] NeurIPS22 69.4 83.0 84.9 76.4 85.2 84.6 73.3 71.1 86.3 79.0 72.1 88.7 79.5 82.6
TPDS-B32 [49] IJCV24 67.7 82.7 86.1 76.5 88.0 86.3 73.6 69.4 87.1 79.3 70.5 89.7 79.7 78.0
MMGA-B32 [7] PR25 68.5 84.6 86.8 78.8 87.5 86.5 75.9 68.3 86.3 80.4 69.9 88.6 80.2 89.2
CausalDA-C-RN 60.1 85.6 86.2 77.2 86.0 86.3 76.6 61.0 86.5 77.5 61.4 86.2 77.6 89.3
CausalDA-C-B32 74.2 90.8 90.3 82.3 91.0 89.9 81.0 74.8 90.0 82.5 74.1 90.9 84.3 90.3
TABLE IV: Closed-set results (%) on DomainNet-126. SF and ViL means source-free and ViL model-required, respectively.
@@
CLIP-RN [42] ICML21 70.2 87.1 65.4 67.9 87.1 65.4 67.9 70.2 65.4 67.9 70.2 87.1 72.7
CLIP-B32 [42] ICML21 76.2 89.0 73.6 77.9 89.0 73.6 77.9 76.2 73.6 77.9 76.2 89.0 79.2
DAPL-RN [14] TNNLS23 72.4 87.6 65.9 72.7 87.6 65.6 73.2 72.4 66.2 73.8 72.9 87.8 74.8
ADCLIP-RN [47] ICCV23W 71.7 88.1 66.0 73.2 86.9 65.2 73.6 73.0 68.4 72.3 74.2 89.3 75.2
SHOT [31] ICML20 63.5 78.2 59.5 67.9 81.3 61.7 67.7 67.6 57.8 70.2 64.0 78.0 68.1
GKD [48] IROS21 61.4 77.4 60.3 69.6 81.4 63.2 68.3 68.4 59.5 71.5 65.2 77.6 68.7
NRC [64] NeurIPS21 62.6 77.1 58.3 62.9 81.3 60.7 64.7 69.4 58.7 69.4 65.8 78.7 67.5
AdaCon [5] CVPR22 60.8 74.8 55.9 62.2 78.3 58.2 63.1 68.1 55.6 67.1 66.0 75.4 65.4
CoWA [26] ICML22 64.6 80.6 60.6 66.2 79.8 60.8 69.0 67.2 60.0 69.0 65.8 79.9 68.6
PLUE [32] CVPR23 59.8 74.0 56.0 61.6 78.5 57.9 61.6 65.9 53.8 67.5 64.3 76.0 64.7
TPDS [49] IJCV24 62.9 77.1 59.8 65.6 79.0 61.5 66.4 67.0 58.2 68.6 64.3 75.3 67.1
SHOT-B32 [31] ICML20 72.8 84.5 72.0 75.1 85.2 72.2 78.2 75.5 71.6 75.2 73.6 83.2 76.6
GKD-B32 [48] IROS21 73.7 84.4 73.6 76.1 85.4 73.6 80.4 76.9 73.9 74.9 74.8 82.1 77.5
NRC-B32 [64] NeurIPS21 72.2 83.4 72.5 75.1 84.8 72.6 78.3 76.7 74.0 75.0 73.7 82.6 76.7
AaD-B32 [63] NeurIPS22 72.2 84.2 72.0 75.3 85.3 72.6 78.2 77.4 73.6 74.8 74.7 82.8 76.9
TPDS-B32 [49] IJCV24 72.5 82.8 71.8 77.7 84.7 72.9 81.9 76.5 72.8 73.5 74.2 83.0 77.0
MMGA-B32 [7] PR25 75.4 86.5 69.9 74.6 88.3 72.3 75.9 76.1 70.9 76.2 75.6 87.4 77.4
CausalDA-C-RN 75.4 88.2 72.0 75.8 88.3 72.1 76.1 75.6 71.2 77.6 75.9 88.2 78.0
CausalDA-C-B32 79.5 89.8 76.9 81.6 90.1 77.5 81.7 80.4 76.6 82.5 80.1 89.9 82.2
@@ 4.3 Closed-set SFDA
Datasets. The evaluation is based on three challenging benchmarks. Office-Home [57] is a middle-scale dataset. It contains 15k images belonging to 65 categories from working or family environments, being divided into four domains, i.e., Artistic, Clip Art, Product, and Real-word. VisDA [40] is a large-scale dataset. It includes 12 types of Synthetic to Real transfer recognition tasks. The source domain contains 152k synthetic images, whilst the target domain has 55k real object images from COCO. DomainNet-126 [39] is another large-scale dataset. As a subset of DomainNet containing 600k images of 345 classes from 6 domains of different image styles, this dataset has 145k images from 126 classes, sampled from 4 domains, Clipart, Painting, Real, Sketch, as [39] identifies severe noisy labels in the dataset.
TABLE V: One step adapting results (%) on Office-Home in Generalized SFDA setting. S, T, H are the results on the source and target domains, and harmonic mean accuracy, respectively; the bracket values in red are the gap between S and T; WAD means With Anti-forgetting Design.
@@
S T H
Source 98.1 59.2 73.1
SHOT [31] ICML20 84.2 71.8 77.1
GKD [48] IROS21 86.8 72.5 78.7
NRC [64] NeurIPS21 91.3 72.3 80.4
AdaCon [5] CVPR22 88.2 65.0 74.4
CoWA [26] ICML22 91.8 72.4 80.6
PLUE [32] CVPR23 96.3 66.9 78.4
TPDS [49] IJCV24 83.8 73.5 78.0
GDA [66] ICCV21 80.0 70.2 74.4
PSAT-RN [50] TMM24 85.3 72.6 78.4
SHOT-B32 [31] ICML20 81.9 79.5 80.7
GKD-B32 [48] IROS21 89.2 79.5 84.1
NRC-B32 [64] NeurIPS21 81.1 78.6 79.8
TPDS-B32 [49] IJCV24 88.9 79.7 84.1
PSAT-B32 [50] TMM24 89.5 80.7 84.9
MMGA-B32 [7] PR25 86.9 80.2 83.4
CausalDA-C-RN 85.0 77.6 80.7
CausalDA-C-B32 86.1 84.3 85.2
TABLE VI: Continual adaptation results (%) on Office-Home under Generalized SFDA setting, the first column of each sub-table indicates the adaptation sequence. $\downarrow$ means the average accuracy drop of a test domain on the adaptation path compared with the performance when the domain is first seen.
CausalDA-C-RN
@@
Competitors. To evaluate CausalDA, we select 14 state-of-the-art methods in three groups. (i) The first is the source model and CLIP that lay the basis for CausalDA. (ii) The second is the UDA method DAPL [14], PADCLIP [25], ADCLIP [47] that introduces prompt learning to boost the cross-domain transfer. (iii) The third includes 9 SFDA methods: SHOT [31], GKD [48], NRC [64], AaD [63], AdaCon [5], CoWA [26], PLUE [32], TPDS [49], and MMGA [7]. By idea, SHOT, GKD, CoWA are pseudo-labels-based methods; NRC, AaD, PLUE, AdaCon exploit the data geometry, e.g., the nearest neighbor, neighborhood, prototypes; TPDS predicts the target probability distribution using error control; MMGA aligns the target domain with the multimodal space represented by CLIP.
For comprehensive comparisons, we implement two variants: (i) CausalDA-C-RN (base) and (ii) CausalDA-C-B32 (premium). The key distinction lies in the backbone of the CLIP image encoder. Specifically, for CausalDA-C-RN, ResNet101 is employed on the VisDA dataset, while ResNet50 is used on the Office-Home and DomainNet-126 datasets. On the other hand, CausalDA-C-B32 adopts ViT-B/32 [15] as the backbone across all datasets. This extends to all other methods using CLIP.
For an SFDA-focused comparison, we integrate the same CLIP model with existing SFDA methods, utilizing the CLIP’s visual backbone (ViT B/32) as the feature extractor, followed by the original training algorithm. As such, all compared methods benefit from the CLIP pre-trained knowledge in a design-generic manner. Specifically, we select five updated SFDA methods as comparisons, including SHOT-B32, GKD-B32, NRC-B32, AaD-B32, and TPDS-B32.
@@
On the other hand, CausalDA-C-RN defeats CLIP-RN on all tasks, achieving the improvement of 4.5%, 5.6%, and 5.3% on Office-Home, VisDA, and DomainNet-126 on average, respectively. Similarly, compared with CLIP-B32, the corresponding improvement of CausalDA-C-B32 are changed to 5.1%, 5.1% and 3.0%. This indicates the effect of the prompt learning to discover the external causal elements.
Among the “-B32” models, we have two observations. First, CausalDA-C-B32 outperforms all conventional methods across all tasks on the three datasets, confirming the effectiveness of our design. Second, on the VisDA dataset, improvements with the B32 versions are limited, and they are even outperformed by the ResNet version, such as AaD and TPDS. This is because the rendered synthetic data (the source domain in VisDA) lacks visual details, appearing primarily in shades of white and gray light and shadow. When processed in ViT, this results in confused semantics. As for AaD and TPDS, contrastive computation (AaD) and chain-like search (TPDS) amplify the noise in the confused semantics, leading to uncontrolled error propagation.
TABLE VII: Partial-set, Open-set results (%) on Office-Home.
@@
SHOT [31] ICML20 79.3 SHOT [31] ICML20 72.8
HCL [20] NeurIPS21 79.6 HCL [20] NeurIPS21 72.6
CoWA [26] ICML22 83.2 CoWA [26] ICML22 73.2
AaD [63] NeurIPS22 79.7 AaD [63] NeurIPS22 71.8
CRS [68] CVPR23 80.6 CRS [68] CVPR23 73.2
SHOT-B32 [31] ICML20 80.6 SHOT-B32 [31] ICML20 73.3
AaD-B32 [63] NeurIPS22 81.2 AaD-B32 [63] NeurIPS22 72.4
CausalDA-C-RN 82.9 CausalDA-C-RN 79.6
CausalDA-C-B32 87.1 CausalDA-C-B32 84.0
4.4 Generalized SFDA
Dataset. We evaluate CausalDA on Office-Home, following previous Generalized SFDA works [66, 50].
Evaluation protocol. Unlike Closed-set SFDA, the Generalized SFDA problem highlights the anti-forgetting ability on the seen source domain. In terms of evaluation rule, the same as [66], we adopt the harmonic mean accuracy that is computed by ${H}=({2*A_{s}*A_{t}})/({A_{s}+A_{t}})$ where $A_{s}$ and $A_{t}$ are the accuracy on the source domain and the target domain, respectively. Note that the $A_{s}$ is computed based on the source-testing set. The same as [66, 50], on the source domain, the ratio of training and testing sets is 9:1.
Competitors. As listed in Tab. V, two Generalized SFDA comparisons, GDA [66] and PSAT [50], are additionally selected. GDA integrated domain attention with a local clustering-based self-supervised learning method. Essentially, knowledge of old tasks is preserved by constraining parameters near their original values. Instead, PSAT enforces the adaptation process to remember the source domain by imposing source guidance, building a target domain-centric anti-forgetting mechanism. As CausalDA, we make two PSAT variants with ResNet50 (PSAT-RN) and ViT (PSAT-B32).
One step adapting. Tab. V reports the results, including source and target domain accuracies and their average. Among the “-B32” methods, CausalDA-C-B32 (without any anti-forgetting design) defeats the best specially designed model PSAT-B32 (with an anti-forgetting mechanism) with a lead of 0.3% in harmonic mean accuracy. CausalDA-C-RN outperforms other alternatives with the same backbone. These results verify the competitiveness of our models in one-step adaptation.
[…]
4.5 Open-set & Partial-set SFDA
Dataset. We adopt Office-Home for evaluation [31, 68].
Competitors. Except Source, CLIP, SHOT, and CoWA as in the Closed-set setting, we consider two more contrastive learning based SFDA methods, HCL [20] and CRS [68].
Comparison results. As reported in Tab. VII, CausalDA-C-B32 achieves the highest average accuracy, outperforming all competing methods by margins of 5.9% and 10.7% over AaD-B32 in the Partial-set setting and SHOT-B32 in the Open-set setting, respectively. Similarly, it significantly outperforms SHOT-B32 and AaD-B32 under both Partial-set and Open-set settings, indicating the efficacy of the proposed design. Only CoWA beats CausalDA-C-RN in the Partial-set setting with a tiny lead 0.3%. The results indicate that CausalDA is competitive for the Partial-set and Open-set settings.